# Large Scale Transfer Learning for Tabular Data via Language Modeling

### v1

@@ Abstract
Tabular data – structured, heterogeneous, spreadsheet-style data with rows and columns – is widely used in practice across many domains. However, while recent foundation models have reduced the need for developing task-specific datasets and predictors in domains such as language modeling and computer vision, this transfer learning paradigm has not had similar impact in the tabular domain. In this work, we seek to narrow this gap and present TabuLa-8B, a language model for tabular prediction. We define a process for extracting a large, high-quality training dataset from the TabLib corpus, proposing methods for tabular data filtering and quality control. Using the resulting dataset, which comprises over 1.6B rows from 3.1M unique tables, we fine-tune a Llama 3-8B large language model (LLM) for tabular data prediction (classification and binned regression) using a novel packing and attention scheme for tabular prediction. Through evaluation across a test suite of 329 datasets, we find that TabuLa-8B has zero-shot accuracy on unseen tables that is over 15 percentage points (pp) higher than random guessing, a feat that is not possible with existing state-of-the-art tabular prediction models (e.g. XGBoost, TabPFN). In the few-shot setting (1-32 shots), without any fine-tuning on the target datasets, TabuLa-8B is 5-15 pp more accurate than XGBoost and TabPFN models that are explicitly trained on equal, or even up to 16$\times$ more data. We release our model, code, and data along with the publication of this paper.
@@ 1 Introduction
Here, we aim to bridge this gap. We introduce TabuLa-8B, a language model for tabular prediction which can flexibly solve classification tasks across unseen domains. Given a small number of examples (shots), without any fine-tuning, TabuLa-8B outperforms state-of-the-art gradient-boosted decision trees and tabular deep learning methods that are explicitly trained on the target data (see Figure 1). Furthermore, TabuLa-8B is capable of zero-shot prediction, a behavior which is not possible using these prior methods. To enable these results, we build a new dataset for tabular prediction, The Tremendous TabLib Trawl (T4), that allows us to scale up training by several orders of magnitude (10,000$\times$ more data) relative to previous work.
@@ 1.1 Our Contributions
Summarizing, this paper has the following three main contributions:
TabuLa-8B, a Tabular Prediction Model: We build TabuLa-8B (Tabular LLama 3 - 8B), a model for prediction on tabular data. On an evaluation suite consisting of 329 tables drawn from five tabular benchmarks, TabuLa-8B has zero-shot accuracy 17 percentage points (pp) above random guessing. In the few-shot setting (1-32 examples), TabuLa-8B is 5-15 pp more accurate than state-of-the-art methods (XGboost, TabPFN) that are trained on equal number of shots, and these methods require 2-8$\times$ more data to achieve the performance of our model.
T4 - A Large Scale, High Quality Training Dataset: We build and release The Tremendous TabLib Trawl (T4), a filtered collection of 3.1M unique tables (consisting of over 1.6B rows, a total of 80B tokens) from TabLib [13]. We detail the recipe used to construct T4, including a suite of methods for filtering web-scale tabular data at several levels (table, row, column), removing unwanted information such as PII and code, and selecting unsupervised prediction targets from these tables.
@@ 3 TabuLa-8B - Model Design and Training
Training Details: The final model is trained for 40k steps with a global batch size of 24 (with sample packing, this is roughly equivalent to a global batch size of 600 rows of tabular data). The model sees roughly 8B tokens during training; we note that is only 10% of the 80Btokens in T4, and less than one one thousandth of TabLib itself. We fully fine-tune all model parameters, as opposed to parameter-efficient fine-tuning, since full-fine tuning consistently benefits from scale [64, 24]. Reproducibility details are given in Appendix B.
@@ 4 Dataset Construction: Building The Tremendous TabLib Trawl (T4)
Final T4 Dataset Summary: Running this entire filtering process (from raw data to serialized examples ready for training) on all 70TB of TabLib using our open-sourced implementation takes about 4 hours on a CPU cluster. It yields a total of 3.1M tables, which equates to a table filtering rate of approximately 98.45%. The resulting dataset contains over 1.6B rows (approximately 80B Llama 3 tokens) for training of the downstream model, and occupies roughly 2TB compressed on disk. We note that 80B tokens is larger than the total number of tokens TabuLa-8B sees during training. Therefore, the model sees each distinct table at most once during training, and our pipeline could be scaled up to support larger models or longer training runs.
@@ 5.3 Baselines
Whenever possible, we perform hyperparameter tuning on XGBoost and TabPFN in order to maximize their performance. See Appendix D.2 for further details on baseline implementation and tuning.
@@ 5.5 Further Robustness Evaluations and Ablation Experiments
Apart from evaluating TabuLa-8B’s transfer learning ability, we also investigate its robustness and the degree to which performance is affected by the order in which columns are presented (serialized); this order invariance is cited as a necessary attribute of tabular foundation models in [58]. We present these experiments in Appendix F.4. Our results demonstrate that changing column order does not alter performance in a statistically significant way, but that there may be a small ($\sim 1$pp) drop which we hypothesize is due to manual ordering of tabular columns in certain benchmark datasets.
Second, we evaluate the impact of the attention masking scheme introduced and described in Section 3. We conduct an ablation study, replacing this component of the model with a sample-wise causal attention (the same form of attention used during standard language model training, where attention across documents is prevented). Our results, detailed in Appendix F.1 and Figure 6, illustrate that this modification is central to the few-shot learning capabilities of TabuLa-8B: when our mechanism is replaced with sample-wise attention the resulting model does not demonstrate few-shot learning capacity, and its performance degrades for $k\geq 16$ (see Figure 6).
@@ 5.6 Assessing the Potential Impact of Data Contamination
Given that T4 consists of 3.1Mtables sourced from public data sources (Common Crawl, Github) and that our evaluations are also comprised of public benchmarks, we investigate the extent and possible impact of data contamination – that is, training datasets that are part of the evaluation suite. In Section G, we explain our methodology to test for the potential presence of benchmark datasets in T4. Using a conservative identification strategy based on column matching (likely to include false positives). We find at most one-third of benchmark tables may occur at least once in the training set. When training large-scale models for transfer learning, it is not always clear a priori what the eventual application domains will be. Therefore, we believe that it is an important research question to understand the extent to which contamination may affect performance, as contamination may be difficult to prevent in some cases. Initial foundation modeling efforts in non-tabular domains adopted a similar approach, and found mixed or no impact from overlap [40, 38].
We evaluate the impact of contamination in our experimental setup by evaluating TabuLa-8B separately on “potentially contaminated” vs. uncontaminated tables. Our results are shown in the bottom right plot of Figure 4, as well as in Figures 11 and 12. Summarizing, we find no clear evidence that contamination affects model performance on the test suite, or that transfer ability is affected by contamination. In fact, as seen in Figure 4, the gap between TabuLa-8B and XGBoost is in fact larger if we restrict evaluation to the benchmark tables which we verify are not in T4. In addition to verifying that our results continue to hold over a diverse set of tables which we know the model did not see in training, it also shows that having some amount of potential contamination did not upwardly bias our estimate of TabuLa-8B’s transfer learning ability. We hypothesize that the observed gap in Figure 4 is due to our conservative duplication procedure being more likely to flag datasets with generic or common column names, which also leads to worse baseline performance on these tasks. We present more comprehensive investigation on the effects of contamination in Appendix G.
@@ 6 Discussion
This work was supported by a Microsoft Grant for Customer Experience Innovation. This work was also in part supported by the NSF AI Institute for Foundations of Machine Learning (IFML, CCF-2019844), Google, Open Philanthropy, and the Allen Institute for AI. JCP was supported by the Harvard Center for Research on Computation and Society.
We are grateful to Foundry [^1] https://www.mlfoundry.com for providing the compute infrastructure used to train TabuLa-8B. Our research also utilized computational resources and services provided by the Hyak computing cluster at the University of Washington, and from Stability AI. The authors also gratefully acknowledge the Gauss Centre for Supercomputing222www.gauss-centre.eu for funding this project by providing computing time on the GCS Supercomputer JUWELS[28] at Julich Supercomputing Centre (JSC). We are particularly appreciative of support from Jenia Jitsev at JSC.
@@ A.1 Table, Column, and Row Filtering Rules
The following tables describe the entire set of heuristics we use as filters. A precise implementation may be found as a part of our software release associated with this paper. Our pipeline involves a langauge identification step; for language detection, we make use of the fasttext library [27]. [^4] https://pypi.org/project/fasttext-langdetect/
@@
Table English Filtering Drop where a langauge ID model score is below a fixed threshold All downstream benchmark datasets are in English
@@
Table Row Count Drop table with fews than $k$ rows Anecdotally, many “very small” tables in TabLib are general web-text tables not useful/suitable for ML.
@@ A.2 Target Column Selection
A.3 Implementation Details
@@ D.2 Baselines
Describe any baseline models and any special techniques or tuning that differ from what we use in the main text.
For the supervised learning baselines (XGBoost, TabPFN), we conduct 10 independent trials, drawing separate training sets of size $k$ shots, for each value of $k$. We use the full remaining dataset as the test set.
Hyperparameter tuning: For each of the 10 independent trials, we tune the hyperparameters of the model. For XGBoost, we conduct 10 iterations of hyperparameter tuning using the HyperOpt hyperparameter optimization library and the hyperparameter grid defined in [16]. For TabPFN, we conduct a full grid search (since there is only a single hyperparameter).
@@ D.2.1 Llama 3 Base Model
We use the pretrained Llama 3 model available on Hugging Face. This model, we do not modify the tokenizer (i.e. by adding special tokens that are used by TabuLa-8B), but the serialized data format is identical to the format seen by our model during training (Figure 2(b)).
@@ Appendix E Detailed Results
Figure 5: 5(a): Results for 32-shot tasks, with extended curves for baselines. 5(b), 5(c): Results for 64- and 128-shot tasks. All models continue to improve as $k$ increases, but the gap between methods may narrow. However, the nature of the datasets that can be used for 64-shot learning with TabuLa-8B are considerably different – fewer features, with shorter, less semantically-meaningful column names – which may downwardly bias the observed performance of TabuLa-8B with large $k$.

### v2

@@ Abstract
Tabular data – structured, heterogeneous, spreadsheet-style data with rows and columns – is widely used in practice across many domains. However, while recent foundation models have reduced the need for developing task-specific datasets and predictors in domains such as language modeling and computer vision, this transfer learning paradigm has not had similar impact in the tabular domain. In this work, we seek to narrow this gap and present TabuLa-8B, a language model for tabular prediction. We define a process for extracting a large, high-quality training dataset from the TabLib corpus, proposing methods for tabular data filtering and quality control. Using the resulting dataset, which comprises over 2.1B rows from over 4M unique tables, we fine-tune a Llama 3-8B large language model (LLM) for tabular data prediction (classification and binned regression) using a novel packing and attention scheme for tabular prediction. Through evaluation across a test suite of 329 datasets, we find that TabuLa-8B has zero-shot accuracy on unseen tables that is over 15 percentage points (pp) higher than random guessing, a feat that is not possible with existing state-of-the-art tabular prediction models (e.g. XGBoost, TabPFN). In the few-shot setting (1-32 shots), without any fine-tuning on the target datasets, TabuLa-8B is 5-15 pp more accurate than XGBoost and TabPFN models that are explicitly trained on equal, or even up to 16$\times$ more data. We release our model, code, and data along with the publication of this paper.  [^1] For links to all code, data, and model, see Section 7.
@@ 1 Introduction
Here, we aim to bridge this gap. We introduce TabuLa-8B, a language model for tabular prediction which can flexibly solve classification tasks across unseen domains, including where data is scarce. Our methodology expands the scope what is possible in these settings, thereby democratizing access to prediction in low-resource contexts and providing state-of-the-art, training-free transfer learning on any tabular data. Since the model only requires a forward pass to perform inference on the target data, it also avoids the privacy or computational considerations that arise in other approaches that require fine-tuning on local, and potentially sensitive, datasets.
In particular, given a small number of examples (shots), and without any fine-tuning on the task, TabuLa-8B outperforms state-of-the-art gradient-boosted decision trees and tabular deep learning methods that are explicitly trained on the target data (see Figure 1). Furthermore, TabuLa-8B is capable of zero-shot prediction, a behavior which is not possible using these prior methods. To enable these results, we build a new dataset for tabular prediction, The Tremendous TabLib Trawl (T4), that allows us to scale up training by several orders of magnitude (10,000$\times$ more data) relative to previous work.
@@ 1.1 Our Contributions
This paper has the following three main contributions:
TabuLa-8B, a Tabular Prediction Model: We build TabuLa-8B (Tabular Llama 3 - 8B), a model for prediction on tabular data. On an evaluation suite consisting of 329 tables drawn from five tabular benchmarks, TabuLa-8B has zero-shot accuracy 17 percentage points (pp) above random guessing. In the few-shot setting (1-32 examples), TabuLa-8B is 5-15 pp more accurate than state-of-the-art methods (XGboost, TabPFN, CatBoost) that are trained on equal number of shots, and these methods require 2-8$\times$ more data to achieve the performance of our model. TabuLa-8B outperforms a variety of strong tabular baselines and even commercial LLMs such as Claude Instant and Claude 3 Sonnet.
T4 - A Large Scale, High Quality Training Dataset: We build and release The Tremendous TabLib Trawl (T4), a filtered collection of 4M unique tables (consisting of over 2.1B rows, a total of 100B tokens) from TabLib [13]. We detail the recipe used to construct T4, including a suite of methods for filtering web-scale tabular data at several levels (table, row, column), removing unwanted information such as PII and code, and selecting unsupervised prediction targets from these tables.
@@ 3 TabuLa-8B - Model Design and Training
Training Details: The final model is trained for 40k steps with a global batch size of 24 (with sample packing, this is roughly equivalent to a global batch size of 600 rows of tabular data). The model sees roughly 8B tokens during training; we note that is less than 10% of the 100Btokens in T4, and less than one one thousandth of TabLib itself. We fully fine-tune all model parameters, as opposed to parameter-efficient fine-tuning, since full-fine tuning consistently benefits from scale [64, 24]. Reproducibility details are given in Appendix B.
@@ 4 Dataset Construction: Building The Tremendous TabLib Trawl (T4)
Final T4 Dataset Summary: Running this entire filtering process (from raw data to serialized examples ready for training) on all 70TB of TabLib using our open-sourced implementation takes about 4 hours on a CPU cluster. It yields a total of 4M tables, which equates to a table filtering rate of approximately 97.91%. Additional descriptive statistics for the dataset are given in Appendix A.3. The resulting dataset contains over 2.1B rows (approximately 100B Llama 3 tokens) for training of the downstream model, and occupies roughly 2TB compressed on disk. We note that 100B tokens is larger than the total number of tokens TabuLa-8B sees during training. Therefore, the model sees each distinct table at most once during training, and our pipeline could be scaled up to support larger models or longer training runs.
@@ 5.3 Baselines
Whenever possible, we perform hyperparameter tuning on XGBoost and TabPFN in order to maximize their performance. See Appendix D.2 for further details on baseline implementation and tuning. We also provide results comparing to additional supervised baseline models, and to commercial LLMs, in Section E.2.
@@ 5.5 Further Robustness Evaluations and Ablation Experiments
Robustness to Column Ordering. Apart from evaluating TabuLa-8B’s transfer learning ability, we also investigate its robustness and the degree to which performance is affected by the order in which columns are presented (serialized); this order invariance is cited as a necessary attribute of tabular foundation models in [58]. We present these experiments in Appendix F.4. Our results demonstrate that changing column order does not alter performance in a statistically significant way, but that there may be a small ($\sim 1$pp) drop which we hypothesize is due to manual ordering of tabular columns in certain benchmark datasets which sometimes reflects a more “natural” ordering.
Robustness to Feature Dropout. Language models may be uniquely susceptible to small changes in the downstream data; for example, the removal of specific features may affect language models’ prediction performance more than traditional supervised methods. We conduct an ablation study to assess the behavior of TabuLa-8B as columns are removed from the test data. We assess removal both in order of descending and ascending importance. The results of these experiments, in Appendix F.3, demonstrate that TabuLa-8B’s performance declines at a similar rate to an XGBoost model trained directly on the subset of features.
Robustness to Column Header Removal. Another potential risk of tabular language models is that, while these models are able to utilize the semantic information in column names, the model may also be overly reliant on the presence of informative column names. In Appendix F.2 we conduct an ablation study to assess this. The results in Appendix F.2 demonstrate that there is a small decline in performance when column headers are removed (replaced with uninformative headers), but that TabuLa-8B still outperforms baselines across all numbers of shots. We believe that this drop in performance is commensurate to the loss in information when column headers are eliminated.
Importance of Row-Causal Tabular Mask. We evaluate the impact of the attention masking scheme introduced and described in Section 3. We conduct an ablation study, replacing this component of the model with a sample-wise causal attention (the same form of attention used during standard language model training, where attention across documents is prevented). Our results, detailed in Appendix F.1 and Figure 11, illustrate that this modification is central to the few-shot learning capabilities of TabuLa-8B: when our mechanism is replaced with sample-wise attention the resulting model does not demonstrate few-shot learning capacity, and its performance degrades for $k\geq 16$ (see Figure 11).
Influence of the Base LLM. We conduct an ablation study of the base LLM to evaluate how TabuLa-8B improves as the base LLM improves. In particular, we rerun our main training pipeline described in Section 3, but using Llama 1 and 2 as the initial language model instead of Llama 3. These results, provided in Section F.6, demonstrate that TabuLa-8B improves along with the performance of the underlying base model. Taken together, these results highlight how the primary contribution of the paper is not the specific model we produce, as much as it is the methodology we present for generating tabular predictors from base language models. As LLMs continue to improve, so will the tabular models that are produced by applying our training methodology to new LLMs.
@@ 5.6 Assessing the Potential Impact of Data Contamination
Given that T4 consists of 4Mtables sourced from public data sources (Common Crawl, Github) and that our evaluations are also comprised of public benchmarks, we investigate the extent and possible impact of data contamination – that is, training datasets that are part of the evaluation suite. In Section G, we explain our methodology to test for the potential presence of benchmark datasets in T4. Using a conservative identification strategy based on column matching (likely to include false positives). We find at most one-third of benchmark tables may occur at least once in the training set. When training large-scale models for transfer learning, it is not always clear a priori what the eventual application domains will be. Therefore, we believe that it is an important research question to understand the extent to which contamination may affect performance, as contamination may be difficult to prevent in some cases. Initial foundation modeling efforts in non-tabular domains adopted a similar approach, and found mixed or no impact from overlap [40, 38].
We evaluate the impact of contamination in our experimental setup by evaluating TabuLa-8B separately on “potentially contaminated” vs. uncontaminated tables. Our results are shown in the bottom right plot of Figure 4, as well as in Figures 17 and 18. Summarizing, we find no clear evidence that contamination affects model performance on the test suite, or that transfer ability is affected by contamination. In fact, as seen in Figure 4, the gap between TabuLa-8B and XGBoost is in fact larger if we restrict evaluation to the benchmark tables which we verify are not in T4. In addition to verifying that our results continue to hold over a diverse set of tables which we know the model did not see in training, it also shows that having some amount of potential contamination did not upwardly bias our estimate of TabuLa-8B’s transfer learning ability. We hypothesize that the observed gap in Figure 4 is due to our conservative duplication procedure being more likely to flag datasets with generic or common column names, which also leads to worse baseline performance on these tasks. We present more comprehensive investigation on the effects of contamination in Appendix G.
@@ 6 Discussion
7 Accessing Open-Source Code, Data, and Model Weights
TabLib Preprocessing Code: A Python module for filtering TabLib, tabliblib, along with scripts and configurations used to perform the filtering, are available at https://github.com/mlfoundations/tabliblib.
Model Training and Inference Code: We provide rtfm, a Python module used to train TabuLa-8B, perform inference and evaluation, and process data, at https://github.com/mlfoundations/rtfm.
T4 Dataset: The T4 dataset is available via public credentialized access on Hugging Face datasets at https://huggingface.co/datasets/mlfoundations/t4-full. Because the dataset is derived from TabLib, users must first obtain permission to access TabLib at https://huggingface.co/datasets/approximatelabs/tablib-v1-full.
Evaluation Datasets: The full evaluation suite used to evaluate TabuLa-8B is available via Hugging Face Datasets at https://huggingface.co/datasets/mlfoundations/tabula-8b-eval-suite. Each dataset includes: a CSV file containing the raw data; a TableShift [16] FeatureList JSON object; and a YAML file with associated metadata.
Model Weights: TabuLa-8B weights are available via Hugging Face at https://huggingface.co/mlfoundations/tabula-8b.
@@ 7 Accessing Open-Source Code, Data, and Model Weights
JG was supported by a Microsoft Grant for Customer Experience Innovation. This work was also in part supported by the NSF AI Institute for Foundations of Machine Learning (IFML, CCF-2019844), Google, Open Philanthropy, and the Allen Institute for AI. JCP was supported by the Harvard Center for Research on Computation and Society.
We are grateful to Foundry [^2] https://www.mlfoundry.com for providing the compute infrastructure used to train TabuLa-8B. Our research also utilized computational resources and services provided by the Hyak computing cluster at the University of Washington, and from Stability AI. The authors also gratefully acknowledge the Gauss Centre for Supercomputing333www.gauss-centre.eu for funding this project by providing computing time on the GCS Supercomputer JUWELS[28] at Julich Supercomputing Centre (JSC). We are particularly appreciative of support from Jenia Jitsev at JSC.
@@ A.1 Table, Column, and Row Filtering Rules
The following tables describe the entire set of heuristics we use as filters. A precise implementation may be found as a part of our software release associated with this paper. Our pipeline involves a language identification step; for language detection, we make use of the fasttext library [27]. [^5] https://pypi.org/project/fasttext-langdetect/
@@
Table English Filtering Drop where a language ID model score is below a fixed threshold All downstream benchmark datasets are in English
@@
Table Row Count Drop table with fewer than $k$ rows Anecdotally, many “very small” tables in TabLib are general web-text tables not useful/suitable for ML.
@@ A.2 Target Column Selection
A.3 Descriptive Statistics
This section provides some basic descriptive statistics of the final dataset, shown in Figures 5 and 6. Figure 5(a) shows that T4 represents a wide variety of data types across its tables, but that data are primarily represented as float, int, and object (string/categorical) data types. Figure 5(b) shows that, while most tables in T4 have little to no missing data, tables can have as much as 10% of data missing (a maximum threshold enforced by heuristics described above).
Figure 6 provides a sense of the “shape” of tables in T4, showing that roughly 30% of tables have 64 columns, and roughly 30% have $1,000$ columns, the minimum and maximum number of rows set by our heuristic filters. In contrast, nearly all tables have fewer than 50 columns, with only a very small fraction ($<0.01\%$) having 500 or more columns (in practice, rows from tables with 500 columns would almost never fit inside the context window of T4 after serialization and tokenization).
Figure 5: Summary metrics for the T4 dataset.
Figure 6: Distribution of counts of rows per table and columns per table in T4.
A.4 Implementation Details
@@ D.2 Baselines
For the supervised learning baselines (XGBoost, TabPFN, CatBoost, Logistic Regression), we conduct 10 independent trials, drawing separate training sets of size $k$ shots, for each value of $k$. We use the full remaining dataset as the test set.
Hyperparameter tuning: For each of the 10 independent trials, we tune the hyperparameters of the model. For XGBoost, we conduct 10 iterations of hyperparameter tuning using the HyperOpt hyperparameter optimization library and the hyperparameter grid defined in [16]. For TabPFN and L2-regularized Logistic Regression, we conduct a full grid search (since there is only a single hyperparameter).
@@ D.2.1 Llama 3 Base Model
We use the pretrained Llama 3 model available on Hugging Face. For this model, we do not modify the tokenizer (i.e. by adding special tokens that are used by TabuLa-8B), but the serialized data format is identical to the format seen by our model during training (Figure 2(b)).
@@ E.1 Results Beyond 32 Shots
Figure 7: 7(a): Results for 32-shot tasks, with extended curves for baselines. 7(b), 7(c): Results for 64- and 128-shot tasks. All models continue to improve as $k$ increases, but the gap between methods may narrow. However, the nature of the datasets that can be used for 64-shot learning with TabuLa-8B are considerably different – fewer features, with shorter, less semantically-meaningful column names – which may downwardly bias the observed performance of TabuLa-8B with large $k$.
E.2 Additional Baseline Comparisons
In this section, we provide comparisons to additional baselines not included in the main text. These include supervised baselines (Logistic Regression, CatBoost), and commercial LLMs (variants of Claude [^11] https://www.anthropic.com/claude).
For the Logistic Regression and Catboost baselines, we follow the same hyperparameter tuning and evaluation procedure described in the main text. These results, along with our main results, are presented in Figure 8.
Figure 8: Few-shot curves for TabuLa-8B along with additional supervised baselines (Logistic Regression, CatBoost).
We conduct additional comparisons to strong commercial LLMs. In particular, we compare TabuLa-8B to two variances of Anthropic’s Claude models: (1) Claude Instant, a fast, relatively small model likely to be on the same order of magnitude of compute as TabuLa-8B (the exact parameter counts of all Claude models are not publicly disclosed); and (2) Claude 3 Sonnet, a highly performant instruction-tuned LLM likely to be larger than TabuLa-8B both in terms of parameter count and total training compute. Due to the cost of accessing these commercial models, we conduct this evaluation on a subset of our evaluation suite.
Figure 9 shows that TabuLa-8B significantly outperforms both models. Figure 9 shows two particularly interesting results: first, both Claude models show little to no improvement with increasing number of shots. This is also consistent with the behavior of the base Llama 3 model. We hypothesize that this behavior demonstrates the value of explicit training for few-shot learning, and likely highlights the gap betweenn more generalized, task-agnostic training of these models relative to the task-specific training of TabuLa-8B. Second, the commercial models perform worse than any other baseline, on average, beyond 3 shots (below which most baselines perform similarly). We hypothesize that this is due, at least in part, to the instruction-following training of the Claude models; no other set of models in our comparison undergo this second post-training phase likely to be utilized in the Claude training pipeline. This instruction-following may improve models alignment with user intentions but could decrease their ability to explicitly learn from data.
Figure 9: Comparison of few-shot performance with commercial models, Claude Instant and Claude 3 Sonnet, along with other baselines.
E.3 Performance on Numeric Tasks
One concern with tabular LLMs is their perceived inability to represent numeric data: since language models represent all data as tokens in a learned embedding space, it may be more challenging for language models to learn the complex relationships between numeric features which are required for many classification tasks.
In order to provide evidence of TabuLa-8B’s performance on numeric data, in this section we report two different slices of our main results. First, we present results on the subset of our evaluation tasks that contain at least one numeric column (int or float data type). This excludes tables that are strictly non-numeric. Second, we present results on only the subset of our evaluation tasks that contain entirely numeric data. We note that both of these subsets exclude tables with purely textual data – precisely the datasets where we might expect language models to perform strongly.
The results on these two subsets of our evaluation suite are shown in Figure 10.
Figure 10(a) shows that, on the evaluation subset where all tables contain at least one numeric column, TabuLa-8B still outperforms baselines across all numbers of shots. Figure 10(a) demonstrates that the mere presence of numeric features does not erode the performance of TabuLa-8B relative to existing SOTA baseline methods.
Figure 10(b) shows that, on the evaluation subset where all tables contain only numeric columns, TabuLa-8B performs on par with existing SOTA baselines but generally only matches their performance. This result is perhaps unsurprising, as numeric-only data is the most advantageous setting for GBDT models and TabPFN (as GBDTs can directly learn splits over numeric values, and TabPFN is exclusively trained on numeric features). However, the ability of TabuLa-8B to match these strong baselines, while exceeding them on non-numeric data and possessing capabilities no other baseline possesses (such as zero-shot prediction and transfer), is an indication of its strength and potential utility as a general tabular classifier.
Figure 10: Few-shot curves for TabuLa-8B on the subset of evaluation tasks that contain at least one numeric column (10(a)) and on the subset of evaluation tasks that contain entirely numeric columns (10(b)).